# NEZHA: A Zero-sacrifice and Hyperspeed Decoding Architecture for Generative Recommendations

### v1

@@ Abstract.
Generative Recommendation (GR), powered by Large Language Models (LLMs), represents a promising new paradigm for industrial recommender systems. However, their practical application is severely hindered by high inference latency, which makes them infeasible for high-throughput, real-time services and limits their overall business impact. While Speculative Decoding (SD) has been proposed to accelerate the autoregressive generation process, existing implementations introduce new bottlenecks: they typically require separate draft models and model-based verifiers, requiring additional training and increasing the latency overhead. In this paper, we address these challenges with NEZHA, a novel architecture that achieves hyperspeed decoding for GR systems without sacrificing recommendation quality. Specifically, NEZHA integrates a nimble autoregressive draft head directly into the primary model, enabling efficient self-drafting. This design, combined with a specialized input prompt structure, preserves the integrity of sequence-to-sequence generation. Furthermore, to tackle the critical problem of hallucination—a major source of performance degradation—we introduce an efficient, model-free verifier based on a hash set. We demonstrate the effectiveness of NEZHA through extensive experiments on public datasets and have successfully deployed the system on Taobao since October 2025, driving the billion-level advertising revenue and serving hundreds of millions of daily active users.
@@
$\dagger$ Work was conducted during the internship of Yejing Wang at Alibaba.
* Xiangyu Zhao is the corresponding author.
copyright: acmcopyrightjournalyear: 2026doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
The extraordinary capabilities of Large Language Models (LLMs) have catalyzed a wide range of applications in recommender systems, which utilizes the LLM abilities including world knowledge integration (Xi et al., 2024a), representation learning (Ren et al., 2024), and semantic understanding (Bao et al., 2025). Among these LLM-driven solutions, Generative Recommendation (GR) (Rajput et al., 2023; Zheng et al., 2024) has rapidly emerged as a dominant paradigm due to its outstanding performance, as well as its notable strengths in addressing the cold-start problem and enhancing recommendation diversity (Rajput et al., 2023). Consequently, GR has seen widespread adoption in various industrial applications and production systems (Zhou et al., 2025; Zheng et al., 2025c; Guo et al., 2025; Wei et al., 2025).
Figure 1. Overview of the GR Deployment Pipeline using the Qwen model as an example. (a) Item Tokenization: An item from the catalog, such as a “phone case for iPhone 17”, is first tokenized into a structured, multi-token semantic ID based on its multi-modal representation. (b) LLM Training: The LLM is then fine-tuned on a large offline dataset ($\mathcal{D}$) of user interactions to learn the patterns for predicting the next item’s semantic ID based on user context. (c) LLM Serving: Finally, the trained model is deployed for real-time serving. Given an incoming user request, it autoregressively generates the semantic ID of the recommended item.
As illustrated in Figure 1, the typical GR deployment involves three stages: (a) item tokenization (Hua et al., 2023; Wang et al., 2024; Zheng et al., 2025b), (b) LLM training (Deng et al., 2025; Yang et al., 2025a), and (c) LLM serving  (Xi et al., 2025; Lin et al., 2025; Xi et al., 2024b). While the first two stages are performed offline, the serving latency of the final stage presents a formidable obstacle to large-scale industrial adoption, particularly in latency-sensitive scenarios. For instance, in Taobao’s search advertising, a business unit that accounts for over a quarter of the company’s revenue, the core service requires a response time of less than 30 milliseconds. However, the latency of the current GR solution we deployed in this scenario exceeds 1 second, thus hindering the full commercial potential of real-time serving.
Algorithm-level solutions for this problem, i.e., efficient LLM inference, broadly fall into two categories (Wan et al., [n. d.]): KV-cache optimization (Xi et al., 2024b) and speculative decoding (SD) (Lin et al., 2025; Zagyva et al., 2025). However, optimizing the KV cache alone is often insufficient: these techniques, e.g., FlashAttention (Dao et al., 2022), yield negligible efficiency gains when generating the extremely short sequences (e.g., 3-4 tokens) common in GR (Dao, [n. d.]). The results presented in Table 1, which already incorporate KV-cache enhancements, still significantly exceed the requirements of industrial applications. Consequently, this paper focuses on developing SD solutions to bridge the efficiency gap. Despite considerable efforts in designing SD frameworks for efficient LLM inference (Zhang et al., 2025; Xia et al., 2024; Zhang et al., 2024), a deployable solution for industrial-scale GR systems remains elusive. Existing SD frameworks tailored for applications exhibit critical shortcomings in both core stages: drafting and verification (Zhang et al., 2024, 2025):
  • Drafting: Current approaches predominantly rely on external draft models, such as smaller language models (Lin et al., 2025; Zagyva et al., 2025) or retrieval models (Ding et al., 2024; Xi et al., 2025). This strategy not only necessitates the costly training and maintenance of an extra model but also introduces additional inference overhead, partially offsetting the intended speed gains.
@@
  • We validate NEZHA’s performance on a large-scale industrial dataset and detail its successful deployment in Taobao, a leading e-commerce platform in China. The system now serves hundreds of millions of daily active users and has yielded substantial advancements in key business.
2. Preliminary
2.1. Generative Recommendation (GR)
The typical process for constructing a GR system is illustrated with the scenario of our E-Commerce Search Advertising in Figure 1, which consists of three steps: item tokenization, LLM training, and LLM serving:
1. Item Tokenization. The first stage involves converting each item into multi-token semantic IDs (Rajput et al., 2023). In this step, quantization methods (e.g., RQ (Lee et al., 2022), PQ (Jegou et al., 2010)) are usually applied to map item contextual representations to discrete codes. These codes are then added to the LLM’s vocabulary as special tokens. For instance, this paper represents an item $v_{i}$ as with a three-token ID $t^{i}_{1},t^{i}_{2},t^{i}_{3}]$ based on multi-modal representations (visual and textual), a process visualized in Figure 1(a) with a phone case.
2. LLM Training. Next, the LLM is trained to understand item semantics and predict user preferences by learning from historical user behavior sequences. The training data consists of tuples $\mathcal{D}=\{(q,x),y\}$, where $q$ is the search query, $x=[v_{1},\dots,v_{N}]$ is the user interaction history, and $y=v_{y}$ is the ground-truth target item ID (e.g., the item the user clicked). All item IDs are flattened into a sequence of tokens. Specifically, the input token sequence is $\text{<BOS>},q,t^{1}_{1},t^{1}_{2},t^{1}_{3},\dots,t^{N}_{1},t^{N}_{2},t^{N}_{3},t^{y}_{1},t^{y}_{2},t^{y}_{3},\text{<EOS>}]$ and the loss is only calculated on $t^{y}_{1},t^{y}_{2},t^{y}_{3}]$. Figure 1(b) visualizes this process with $N=1$.
3. LLM Serving. Finally, the trained model is deployed to generate recommendations for live user requests (Figure 1(c)). For an incoming search query $q$ from a user with history $x$, the model autoregressively generates the semantic ID of the predicted item, token by token, usually with Beam Search (Freitag and Al-Onaizan, 2017). Denoting the LLM as $\mathcal{M}_{p}$, this inference process is formally described in Algorithm 1. The algorithm unifies the beam search steps by initializing the candidate set with a single empty element ($\hat{Y}_{0}=\{\varnothing\}$ in line 1)). Each step then iterates through the current set of top beams (line 3). For each beam, it generates new, longer sequences by prefilling the context and predicting the next possible tokens (lines 4-7). After all candidates have been expanded, the algorithm prunes the resulting set, retaining only the top $K$ candidates with the highest cumulative probabilities (line 9) for the next iteration or output (line 10&12).
A critical bottleneck hinders the fully deployment of GR in high-throughput businesses with strict latency requirements. The issue stems from the autoregressive decoding process in beam search, where the LLM $\mathcal{M}_{p}$ is called iteratively to prefill the context for each new candidate sequence (line 4 in Algorithm 1). Specifically, the LLM is invoked $(K\times(L-1))$ times after the initial context prefill, where $K$ is the beam size and $L$ is the number of tokens in the semantic ID. While $L$ is configurable for different GR systems, $K$ is inevitably large in specific applications; for instance, in item retrieval scenarios, $K$ can be as large as 1,000. Although standard optimizations like KV-Caching (Xi et al., 2024b; Dao et al., 2022) can accelerate these LLM calls, they remain insufficient as evident in Table 1, which shows that even with KV-Cache, the beam search decoding phase requires nearly triple the time of the initial prefilling for 0.6B-parameterized LLM. Consequently, developing novel acceleration techniques is imperative for the industrial adoption of LLM-based GR systems.
@@ 2.2. Speculative Decoding (SD)
Beyond KV-Caching, SD has emerged as the primary choice for algorithm-level acceleration technique (Xia et al., 2024; Zhang et al., 2025). It has garnered significant attention in academia (Lin et al., 2025; Cai et al., 2024; Li et al., 2024) and recently seen successful deployments in industrial applications (Xi et al., 2025; Zagyva et al., 2025).
We illustrate the SD for GR in Algorithm 2, where we configure the draft model to generate the entire multi-token semantic ID in a single step. For clarity, we denote the Algorithm 1 as $\mathrm{BeamSearch}(\cdot)$ in Algorithm 2, which alternatively predicts future tokens with the draft model $\mathcal{M}_{q}$ based on the given initialization $\hat{Y}_{l-1}$ and returns the intermediate candidate set $\hat{Y}_{l}$ for any step $l$, not just the final result $\hat{Y}_{L}$. The core process involves two main steps: drafting and verification. In the drafting stage, SD employs a smaller, faster draft model $\mathcal{M}_{q}$ to generate a sequence of candidate tokens for the larger, more powerful target LLM $\mathcal{M}_{p}$ (line 3). For example, a 68M draft model might be used to generate outputs for a 7B target model, as in AtSpeed (Lin et al., 2025). The drafted tokens are then verified using a single, parallel forward pass of the target LLM (line 4). Based on this verification, if the draft matches the target model’s predictions, it is accepted (line 10). If a mismatch is found, the draft is rejected from the point of the first incorrect token (line 5), and a new token is sampled from the target model’s output before the drafting process repeats (line 6-7).
@@ 3.2. Nimble Drafting
Our design for NEZHA directly addresses this structural constraint. By prompting the input with positional placeholder tokens and utilizing an autoregressive draft head, we explicitly provide the model with the positional awareness and sequential modeling capabilities necessary to generate valid, structured IDs.
Specifically, NEZHA first modifies the input prompt for each training instance $\{(q,x),y\}$. Instead of appending the ground-truth semantic ID $y=[t^{y}_{1},t^{y}_{2},t^{y}_{3}]$, we append a sequence of $L$ special tokens as the placeholder for item ID: $\text{<SP\_1>},\dots,\text{<SP\_L>}]$. This is visually revealed by the modification from Figure 1(b) to Figure 2.
The purpose of these placeholders is twofold. First, drawing inspiration from (Samragh et al., 2025), they explicitly encode positional information, prompting the model to leverage its full contextual understanding for each step of the subsequent autoregressive generation. Second, they allow the LLM to pre-compute the hidden states for all future token positions in a single forward pass, as visualized in the cyan block of Figure 2.
@@
Algorithm 3 NEZHA.
@@ 3.3. Efficient Verification
We elaborate the model-free verification for NEZHA in this section.
@@
The highly structured nature of semantic IDs, which are defined by strict combinatorial rules rather than linguistic grammar, enables a far more efficient verification process. This key property allows us to verify a candidate sequence’s validity by performing a near-instantaneous lookup against a pre-computed set of all valid item IDs, instead of executing a costly forward pass. The power of this approach stems from the extreme sparsity of the valid ID space. As supported by both public datasets (Wang et al., 2024) and our own industrial practice, the ratio of valid IDs is exceptionally low (around 0.1% in our system).
To implement this model-free verification with minimal computational overhead, we first encode each multi-token semantic ID into a single, unique integer (Zheng et al., 2025a):
@@
where this equation maps the semantic ID $t_{1}^{i},\dots,t_{L}^{i}]$ to a unique integer index $V_{i}$ using a mixed-radix conversion. $T_{j}$ is the vocabulary size for the $j$-th token position. The product term $\prod_{j=1}^{l-1}T_{j}$ acts as a positional multiplier. For the first token ($l=1$), this product is empty and evaluates to 1 by convention to correctly initialize the formula.
@@
where $\mathcal{V}_{L}$ is the set of integer indices corresponding to the candidate items in the final step $\hat{Y}_{L}$. This set is then filtered in-place according to the verification process detailed in Equation (5), retaining only the indices of valid items, i.e., with indecies included in $\mathcal{V}\cap\mathcal{V}_{L}$. The visualized example in Figure 2 illustrates this process, where $(0,346,122),(12,1,33)$ is included by the updated $\hat{Y}_{L}$ while excluding other candidates.
The impact of this model-free verification is substantial. On our production data, it boosts the ratio of valid drafted candidates from just 43% to over 93%. This directly translates to a 12 percentage point uplift in key offline evaluation metrics. A detailed component-wise analysis with public datasets confirming this contribution is provided in Section 4.3.

### v2

@@ Abstract.
Generative Recommendation (GR), powered by Large Language Models (LLMs), represents a promising new paradigm for industrial recommender systems. However, their practical application is severely hindered by high inference latency, making them infeasible for high-throughput, real-time services and limiting their overall business impact. While Speculative Decoding (SD) has been proposed to accelerate the autoregressive generation process, existing implementations introduce new bottlenecks: they typically require separate draft models and model-based verifiers, which require additional training and increase latency overhead. In this paper, we address these challenges with NEZHA, a novel architecture that achieves hyperspeed decoding for GR systems without sacrificing recommendation quality. Specifically, NEZHA integrates a nimble autoregressive draft head directly into the primary model, enabling efficient self-drafting. This design, combined with a specialized input prompt structure, preserves the integrity of sequence-to-sequence generation. Furthermore, to tackle the critical problem of hallucination—a major source of performance degradation—we introduce an efficient, model-free verifier based on a hash set. We demonstrate the effectiveness of NEZHA through extensive experiments on public datasets and have successfully deployed the system on Taobao since October 2025, achieving 1.2% business improvement, translating to billion-level advertising revenue and serving hundreds of millions of daily active users. The code is available at https://github.com/Applied-Machine-Learning-Lab/WWW2026_NEZHA.
@@
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792797isbn: 979-8-4007-2307-0/2026/04ccs: Applied computing Electronic commerceccs: Information systems Information retrieval
@@ 1. Introduction
The extraordinary capabilities of Large Language Models (LLMs) have catalyzed a wide range of applications in recommender systems that leverage LLMs’ abilities (Zhu et al., 2025; Zhang et al., 2025c; Zhai et al., 2024), including world knowledge integration (Xi et al., 2024a; Fu et al., 2025b), representation learning (Ren et al., 2024), and semantic understanding (Bao et al., 2025; Zhang et al., 2025a). Among these LLM-driven solutions, Generative Recommendation (GR) (Zheng et al., 2024; Ju et al., 2025; Yang et al., 2024; Gao et al., 2025) has rapidly emerged as a dominant paradigm due to its outstanding performance and notable strengths in addressing the cold-start problem and enhancing recommendation diversity (Rajput et al., 2023; Wang et al., 2025b). Consequently, GR has seen widespread adoption in various industrial applications and production systems (Zhou et al., 2025; Zheng et al., 2025c; Guo et al., 2025; Wei et al., 2025; Chen et al., 2025).
@@
The typical GR deployment involves three stages (Li et al., 2025): (a) item tokenization (Hua et al., 2023; Wang et al., 2024; Zheng et al., 2025a), (b) LLM training (Deng et al., 2025; Yang et al., 2025b), and (c) LLM serving  (Xi et al., 2025; Lin et al., 2025; Xi et al., 2024b). While the first two stages are performed offline, the serving latency of the final stage presents a formidable obstacle to large-scale industrial adoption, particularly in latency-sensitive scenarios (Wang et al., 2025a). For instance, in Taobao’s search advertising, a business unit that accounts for over a quarter of the company’s revenue, the core service requires a response time of less than 30 milliseconds. However, the latency of the current GR solution we deployed in this scenario exceeds 1 second, thus hindering the full commercial potential of real-time serving.
Algorithm-level solutions for this problem, i.e., efficient LLM inference, broadly fall into two categories (Wan et al., ): KV-cache optimization (Xi et al., 2024b) and speculative decoding (SD) (Lin et al., 2025; Zagyva et al., 2025). However, optimizing the KV cache alone is often insufficient: these techniques, e.g., FlashAttention (Dao et al., 2022), yield negligible efficiency gains when generating the extremely short sequences (e.g., 3-4 tokens) common in GR (Dao, ). The results presented in Table 1, which already incorporate KV-cache enhancements, still significantly exceed the requirements of industrial applications. Consequently, this paper focuses on developing SD solutions to bridge the efficiency gap. Despite considerable efforts for efficient LLM inference (Zhang et al., 2025b; Xia et al., 2024; Zhang et al., 2024), a deployable solution for industrial-scale GR systems remains elusive. Existing SD frameworks tailored for applications exhibit critical shortcomings in both core stages: drafting and verification (Zhang et al., 2024, 2025b):
  • Drafting: Current approaches predominantly rely on external draft models, such as smaller language models (Lin et al., 2025; Zagyva et al., 2025) or retrieval models (Ding et al., 2024; Xi et al., 2025). This strategy not only necessitates the costly maintenance of an extra model but also introduces additional inference overhead, offsetting the intended speed gains.
@@
  • We validate NEZHA’s performance on a large-scale industrial dataset and detail its successful deployment in Taobao, a leading e-commerce platform in China. The system now serves hundreds of millions of daily active users and drives billion-level increases in advertising revenue.
Significance. Our work establishes a pivotal advancement in deploying generative recommendation (GR) systems under stringent latency constraints, unlocking their practical viability in real-time, large-scale industrial applications. By effectively addressing the critical challenge of excessive response time that has long hindered GR adoption, we not only demonstrate a highly efficient serving framework tailored for time-sensitive scenarios but also validate its substantial business impact through successful deployment in production. This breakthrough redefines the operational boundaries of GR, transforming it from a latency-prohibitive paradigm into a scalable, high-performance solution.
@@
2. Preliminary
2.1. Generative Recommendation (GR)
The typical process for constructing a GR system consists of three steps: item tokenization, LLM training, and LLM serving:
1. Item Tokenization. The first stage involves converting each item into multi-token semantic IDs (Rajput et al., 2023; Xu et al., 2025; Fu et al., 2025a). In this step, quantization methods (e.g., RQ (Lee et al., 2022), PQ (Jegou et al., 2010)) are usually applied to map item contextual representations to discrete codes. These codes are then added to the LLM’s vocabulary as special tokens. In this paper, we represent an item $v_{i}$ as a three-token ID $t^{i}_{1},t^{i}_{2},t^{i}_{3}]$.
2. LLM Training. Next, the LLM is trained to understand item semantics and predict user preferences by learning from historical user behavior sequences. The training data consists of tuples $\mathcal{D}=\{(q,x),y\}$, where $q$ is the search query, $x=[v_{1},\dots,v_{N}]$ is the user interaction history, and $y=v_{y}$ is the ground-truth target item ID (e.g., the item the user clicked). All item IDs are flattened into a sequence of tokens. Specifically, the input token sequence is $\text{<BOS>},q,t^{1}_{1},t^{1}_{2},t^{1}_{3},\dots,t^{N}_{1},t^{N}_{2},t^{N}_{3},t^{y}_{1},t^{y}_{2},t^{y}_{3},\text{<EOS>}]$ and the loss is only calculated on $t^{y}_{1},t^{y}_{2},t^{y}_{3}]$.
3. LLM Serving. Finally, the trained model is deployed to generate recommendations for live user requests. For an incoming search query $q$ from a user with history $x$, the model autoregressively generates the semantic ID of the predicted item, token by token, usually with Beam Search (Freitag and Al-Onaizan, 2017). Denoting the LLM as $\mathcal{M}_{p}$, this inference process is formally described in Algorithm 1. The algorithm unifies the beam search steps by initializing the candidate set with a single empty element ($\hat{Y}_{0}=\{\varnothing\}$ in line 1)). Each step then iterates through the current set of top beams (line 3). For each beam, it generates new, longer sequences by prefilling the context and predicting the next possible tokens (lines 4-7). After all candidates have been expanded, the algorithm prunes the resulting set, retaining only the top $K$ candidates with the highest cumulative probabilities (line 9) for the next iteration or output (line 10&12).
A critical bottleneck hinders the full deployment of GR in high-throughput businesses with strict latency requirements. The issue stems from the autoregressive decoding process in beam search, where the LLM $\mathcal{M}_{p}$ is called iteratively to prefill the context for each new candidate sequence (line 4 in Algorithm 1). Specifically, the LLM is invoked $(K\times(L-1))$ times after the initial context prefill, where $K$ is the beam size and $L$ is the number of tokens in the semantic ID. While $L$ is configurable for different GR systems, $K$ is inevitably large in specific applications; for instance, in item retrieval scenarios, $K$ can be as large as 1,000. Although optimizations like KV-Caching (Xi et al., 2024b; Dao et al., 2022) can accelerate these LLM calls, they remain insufficient, as evident in Table 1, which shows that even with KV-Cache, the beam search decoding phase requires nearly triple the time of the initial prefilling for a 0.6B-parameterized LLM. Consequently, developing novel acceleration techniques is imperative for the industrial adoption of current GR systems.
@@ 2.2. Speculative Decoding (SD)
Beyond KV-Caching, SD has emerged as the primary choice for algorithm-level acceleration techniques (Xia et al., 2024; Zhang et al., 2025b). It has garnered significant attention in academia (Lin et al., 2025; Cai et al., 2024; Li et al., 2024) and recently seen successful deployments in industrial applications (Xi et al., 2025; Zagyva et al., 2025).
We illustrate the SD for GR in Algorithm 2, where we configure the draft model to generate the entire multi-token semantic ID in a single step. For clarity, we denote the Algorithm 1 as $\mathrm{BeamSearch}(\cdot)$ in Algorithm 2, which alternatively predicts future tokens with the draft model $\mathcal{M}_{q}$ based on the given initialization $\hat{Y}_{l-1}$ and returns the intermediate candidate set $\hat{Y}_{l}$ for any step $l$, not just the final result $\hat{Y}_{L}$. The core process involves two main steps: drafting and verification. In the drafting stage, SD employs a smaller, faster draft model $\mathcal{M}_{q}$ to generate a sequence of candidate tokens for the larger, stronger target LLM $\mathcal{M}_{p}$ (line 3). The drafted tokens are then verified using a single, parallel forward pass of the target LLM (line 4). Based on this verification, if the draft matches the target model’s predictions, it is accepted (line 10). If a mismatch is found, the draft is rejected from the point of the first incorrect token (line 5), and a new token is sampled from the target model’s output before the drafting process repeats (line 6-7).
@@ 3.2. Nimble Drafting
Our design for NEZHA directly addresses this structural constraint. By prompting the input with positional placeholder tokens and utilizing an autoregressive draft head, we explicitly provide the model with the positional awareness and sequential modeling capabilities necessary to generate valid, structured IDs. Specifically, NEZHA first modifies the input prompt for each training instance $\{(q,x),y\}$. Instead of appending the ground-truth semantic ID $y=[t^{y}_{1},t^{y}_{2},t^{y}_{3}]$, we append a sequence of $L$ special tokens as the placeholder for item ID: $\text{<SP\_1>},\dots,\text{<SP\_L>}]$.
The purpose of these placeholders is twofold. First, drawing inspiration from (Samragh et al., 2025), they explicitly encode positional information, prompting the model to leverage its full contexts for each step of the autoregressive generation. Second, they allow the LLM to pre-compute the hidden states for all future token positions in a single forward pass, as visualized in the cyan block of Figure 1.
@@
Algorithm 3 The NEZHA Pipeline: Training and Inference.
@@ 3.3. Efficient Verification
We elaborate on the model-free verification in this section.
@@
The highly structured nature of semantic IDs, which are defined by strict combinatorial rules rather than linguistic grammar, enables a far more efficient verification process. This key property allows us to verify a candidate sequence’s validity by performing a near-instantaneous lookup against a precomputed set of all valid item IDs, rather than executing a costly forward pass. The power of this approach stems from the extreme sparsity of the valid ID space. As supported by both public datasets (Wang et al., 2024) and our own industrial practice, the ratio of valid IDs is exceptionally low (around 0.1% in our system). This means our simple verification step can correctly filter out over 99.9% of all drafted candidates.
To implement this model-free verification with minimal computational overhead, we first encode each multi-token semantic ID into a single, unique integer (Zheng et al., 2025b):
@@
where this equation maps the semantic ID $t_{1}^{i},\dots,t_{L}^{i}]$ to a unique integer index $V_{i}$ using a mixed-radix conversion. $T_{j}$ is the vocabulary size for the $j$-th token position. The product term $\prod_{j=1}^{l-1}T_{j}$ acts as a positional multiplier. For the first token ($l=1$), this product is empty and set to 1 by convention to correctly initialize the formula.
@@
where $\mathcal{V}_{L}$ is the set of integer indices corresponding to the candidate items in the final step $\hat{Y}_{L}$. This set is then filtered in-place according to the verification process detailed in Equation (5), retaining only the indices of valid items, i.e., with indices included in $\mathcal{V}\cap\mathcal{V}_{L}$. The visualized example in Figure 1 illustrates this process, where $(0,346,122),(12,1,33)$ is included by the updated $\hat{Y}_{L}$ while excluding other candidates.
The impact of this model-free verification is substantial. On our production data, it boosts the ratio of valid drafted candidates from just 43% to over 93%. This directly translates into a 12 percentage-point uplift in key offline evaluation metrics. A detailed component-wise analysis using public datasets to confirm this contribution is provided in Section 4.3.